OpenStudy (anonymous):

x/3-y/5=1 graph using x & Y intercepts

OpenStudy (anonymous):

x/3-y/5=1 For x-intercept put y=0 Therefore x/3-0=1 i.e. x=3 i.e. (3,0) For y-intercept put x=0 Therefore 0-y/5=1 i.e. y=-5 i.e. (0,-5) Thus graph of the given eq will pass through the points (3,0) and (0,-5).

OpenStudy (anonymous):

Great!! , thanks